OpenStudy (anonymous):

(a,2b),(3a,-b) Find the midpoint of the line segment.

OpenStudy (atlas):

x co-ordinate of the midpoint will be halfway throught the x coordinates of the given points: x = (a+3a)/2 =2a Similarly for y co-ordinate y = (2b-b)/2 =b/2

OpenStudy (anonymous):

.... I think i get it...